# Global analysis of fragmentation functions to charged hadrons with high-precision data from the LHC

### v1

@@ Appendix B Dataset subtraction for $K^{\pm}$ and $p/\bar{p}$
In this section, we continue the description of the effects of data subtraction on FFs. The comparison between alternative fits and the baseline fit for kaons is depicted in Figs. 34-36, and for protons in Figs. 37-39. The figures, labels, observables, etc., are organized similarly to those for pions. For subtractions of datasets from $pp$ collisions, FFs to $K^{+}$ closely align with the baseline fit except for FFs from gluons. After removal of the ALICE 13 TeV dataset, the gluon FF goes slightly outside the error band for $0.50.2<z<$ 0.5. In the large $z$ region ($>0.4$), after removal of the ATLAS 13 TeV dijets data the gluon FF increases considerably similar to that for pions, since they contain most data points and provides dominant constraints for gluon FF.
@@ Appendix C Additional comparison to data
Figure 40: Comparison of predictions from various fragmentation functions (FFs) for $\pi^{\pm}$ and experimental data at ALICE with a center of mass energy of 2.76 TeV. The theoretical results are renormalized by a factor to align with the experimental data in the last bin. In the lower panel, the ratio to experimental data is presented. The blue bands represent experimental errors, while the light and dark green bands indicate scale uncertainties and Hessian uncertainties, respectively.
@@ D.2 Comparison to analytic results
The partonic coefficient functions have been calculated at NLO in QCD Altarelli et al. (1979); Nason and Webber (1994); Furmanski and Petronzio (1982); Graudenz (1994); de Florian et al. (1998); de Florian and Rotstein Habarnau (2013), and at NNLO in QCD recently Goyal et al. (2023); Bonino et al. (2024). We find very good agreement between the FMNLO results and the analytic predictions as shown in Fig. 44, and the NLO corrections are around 10-40%.

### v2

@@ Appendix B Dataset subtraction for $K^{\pm}$ and $p/\bar{p}$
In this section, we continue the description of the effects of data subtraction on FFs. The comparison between alternative fits and the baseline fit for kaons is depicted in Figs. 37-39, and for protons in Figs. 40-42. The figures, labels, observables, etc., are organized similarly to those for pions. For subtractions of datasets from $pp$ collisions, FFs to $K^{+}$ closely align with the baseline fit except for FFs from gluons. After removal of the ALICE 13 TeV dataset, the gluon FF goes slightly outside the error band for $0.50.2<z<$ 0.5. In the large $z$ region ($>0.4$), after removal of the ATLAS 13 TeV dijets data the gluon FF increases considerably similar to that for pions, since they contain most data points and provides dominant constraints for gluon FF.
@@ Appendix C Additional comparison to data
Figure 43: Comparison of predictions from various fragmentation functions (FFs) for $\pi^{\pm}$ and experimental data at ALICE with a center of mass energy of 2.76 TeV. The theoretical results are renormalized by a factor to align with the experimental data in the last bin. In the lower panel, the ratio to experimental data is presented. The green bars represent experimental errors, while the light and dark green bands indicate scale uncertainties and Hessian uncertainties, respectively.
@@ D.2 Comparison to analytic results
The partonic coefficient functions have been calculated at NLO in QCD Altarelli et al. (1979); Nason and Webber (1994); Furmanski and Petronzio (1982); Graudenz (1994); de Florian et al. (1998); de Florian and Rotstein Habarnau (2013), and at NNLO in QCD recently Goyal et al. (2023); Bonino et al. (2024b). We find very good agreement between the FMNLO results and the analytic predictions as shown in Fig. 47, and the NLO corrections are around 10-40%.